OpenStudy (anonymous):

if x^2-y^2=100 and x-y=20,what is the value of x+y?

OpenStudy (anonymous):

\[x ^{2}-y ^{2}=100\] \[x-y=20\] \[x+y=?\]

OpenStudy (anonymous):

(x+y)(x-y)=x^2-y^2 do you get this part?

OpenStudy (anonymous):

@ineedhelpnow08

OpenStudy (anonymous):

no....

OpenStudy (anonymous):

well if you have an expression in the form of a^2-b^2, then you can factorise it as (a+b)(a-b)

OpenStudy (anonymous):

(a+b)(a-b) a^2+ab-ab-b^2 a^2-b^2

OpenStudy (anonymous):

do you get it now?

OpenStudy (anonymous):

5

OpenStudy (anonymous):

oh..

OpenStudy (anonymous):

so now we can replace x^2-y^2 with (x+y)(x-y) to get (x+y)(x-y)=100

OpenStudy (anonymous):

so like its (20)(x+y)=100?

OpenStudy (anonymous):

since we know (x-y) = 20, we can write (x+y)(20)=100 (x+y)=100/20 x+y=5

OpenStudy (anonymous):

do you get it now?

OpenStudy (anonymous):

thanks!!

OpenStudy (anonymous):

haha... good luck with these type of question..

OpenStudy (anonymous):

lol S.A.T are hard

OpenStudy (anonymous):

haha.. you just gotta work harder...